OpenStudy (anonymous):

Solve the following system of equations. -2x-3y=-9 3x+4y=12 x=? y=?

OpenStudy (anonymous):

x=0 y=3

OpenStudy (anonymous):

3(-2x-3y=-9) 2(3x+4y=12) -6x-9y=-27 6x+8y=24 (cross out -6x +6x) -y=-3 y=3 3x+4y=12 3x+4(3)=12 3x+12=12 3x=0 x=0